OpenStudy (anonymous):

Find the distance between the points (2, -3) and (5, -4). can you please show the steps as well....

OpenStudy (anonymous):

\[\sqrt{10}\]

OpenStudy (anonymous):

apply the distance formula. \[D = \sqrt{(y_2-y_1)^2+(x_2-x_1)^2}\]

OpenStudy (anonymous):

kira is wrong, i guess...........

OpenStudy (anonymous):

\[D = \sqrt{(-4+3)^2+(5-2)^2} = \sqrt{1 + 9} = \sqrt{10}\]

OpenStudy (kira_yamato):

Sorry my fault... Errata:

OpenStudy (anonymous):

correct